# Towards Practical Benchmarking of Data Cleaning Techniques: On Generating Authentic Errors via Large Language Models

### v2

@@ 1. Introduction
Ensuring high data quality is a fundamental challenge in data-driven systems, as erroneous tabular data can terribly degrade the reliability of analytical processes and machine learning models (rahm2000data; chu2016data). To tackle these errors, researchers spend a lot of time designing various data cleaning techniques (chu2013holistic; mahdavi2019raha; rezig2021horizon). As we know, sufficient training data with labeled errors are important for the performance of data cleaning models (rekatsinas2017holoclean). Unfortunately, the manual labeling is very expensive, and obtaining a sufficiently diverse set of real-world errors remains challenging (krishnan2017boostclean). In this sense, it is natural to come up with the idea that we can manually generate errors for the model training.
Existing error generation methods primarily rely on rule-based approaches, with BART (arocena2015messing) being the most representative framework. BART introduces errors into clean datasets by applying predefined constraints, such as functional dependencies (FDs) (mandros2017discovering; fan2008conditional) and denial constraints (DCs) (bian2024discovering; pena2022fast; chu2013discovering), ensuring that the generated errors are systematically detectable. By enforcing a controlled corruption process, BART provides a benchmark for evaluating data cleaning algorithms.
Figure 1. An example of the generated errors by BART (arocena2015messing) and our TableEG, over the Movie dataset.
@@
Second, BART focuses predominantly on constraint-violating errors. Although BART is designed to support a variety of random errors, including typos, duplicated values, outliers, and even missing or bogus values, its rule-based mechanism is inherently limited. In practice, BART struggles to generate missing values (krishnan2017boostclean) that faithfully reflect the nuanced patterns present in real-world datasets and to produce domain-specific semantic inconsistencies (suchanek2011paris). For instance, in the Flights (rekatsinas2017holoclean) dataset, the attribute act_arr_time may contain the value “Not available,” indicating missing or ambiguous data rather than a simple constraint violation. However, BART cannot generate such errors as it assumes all errors stem from various constraint violations. This limitation also weakens its suitability for benchmarking error detection algorithms, which must handle a broader and more diverse range of real-world data corruptions.
@@
Recent advancements in Large Language Models (LLMs) have achieved remarkable success in natural language processing (NLP) tasks, drawing increasing attention to their potential for structured tabular data processing tasks (li2024table; wu2024tablebench; zhang2023tablellama). In this sense, it is intuitive to leverage LLMs’ robust semantic comprehension and expansive knowledge to systematically generate authentic errors. However, unlike natural language, relational data requires maintaining two-dimensional structural integrity while accounting for inter-cell dependencies (e.g., primary-foreign key relationships) and domain-specific constraints (e.g., financial transaction rules). This raises the fundamental questions: Can LLMs generate contextually meaningful errors rather than entirely random or nonsensical corruptions? Do these synthetic errors faithfully replicate the diverse real-world errors so that they can be reliably detected by current error detection algorithms?
@@ 2.1. Types of Data Errors
In data cleaning and quality research, errors are typically defined as deviations of data values from their ground truth (abedjan2016detecting).
@@
These errors can arise from various reasons—such as noise in data acquisition, manual entry mistakes, rule violations, or inconsistencies introduced during data integration. Several studies have proposed classification schemes to categorize these errors either based on their origins (rahm2000data; kim2003taxonomy; hellerstein2013quantitative) or on the violations of constraints and patterns observed in structured data (ilyas2015trends).
Our classification scheme is broadly consistent with the taxonomy proposed by Abedjan et al. (abedjan2016detecting), which, although not exhaustive, serves as a well-established framework for systematically categorizing errors commonly identified in real-world datasets. However, our classification differs in several key aspects. First, we explicitly separate missing values as an independent category, acknowledging their prevalence in real-world data. Second, we incorporate duplicates within the broader category of rule violations. Lastly, our definition of rule violations extends to any value that violates a form of denial constraints, encompassing a wider range of integrity violations. Consequently, we categorize errors into four primary types—outliers, missing values, rule violations, and pattern violations—capturing the most prevalent and impactful error types observed in real-world tabular datasets.
@@
Rule Violations: We employ a set of predefined integrity constraints, primarily in the form of DCs (chu2013discovering)), FDs (mandros2017discovering) and CFDs (fan2008conditional) to safeguard data consistency. Any record that fails to meet these constraints is classified as a rule violation. Such violations often involve contradictory, prohibited, or otherwise infeasible values, thereby severely undermining data integrity. They typically arise from data entry errors, integration inconsistencies, or improper transformations, and are commonly detected through constraint validation or rule-based inference.
@@ 2.2.1. Recent Advancements in Large Language Models (LLMs)
Recent advancements in Large Language Models (LLMs) have demonstrated remarkable capabilities in natural language understanding and generation, primarily trained on vast corpora of sequential text (vaswani2017attention; brown2020language). However, directly applying LLMs to relational data and tabular tasks remains nontrivial due to the inherent structural differences between natural language and relational tables. Unlike textual data, which follows a linear sequence, tabular data is inherently two-dimensional, often governed by schema constraints, cross-column dependencies, and entity relationships (dong2024large). These structural properties impose significant challenges when leveraging LLMs for table-based tasks such as error generation, detection, and repair.
@@ 2.2.2. Challenges of Applying LLMs to Tabular Data
First, traditional LLMs primarily operate under a token-based autoregressive or masked prediction framework, which excels in modeling linguistic dependencies but struggles with structured tabular constraints (yin2020tabert). In relational tables, data integrity often depends on multi-column dependencies, such as functional dependencies (FDs) or domain constraints, which LLMs do not inherently recognize (chu2016data). Consequently, naive applications of LLMs to table-based tasks often result in syntactically plausible yet semantically invalid outputs, where column dependencies are violated, inconsistent values are introduced, or key constraints are broken. Second, LLMs tend to generate errors based on linguistic priors rather than data-specific corruption patterns. Unlike real-world data errors—which may arise from user input mistakes, OCR misrecognition, or systematic extraction faults—LLM-generated corruptions often exhibit artificial substitution biases that fail to reflect natural error distributions. This discrepancy limits the effectiveness of LLM-generated datasets for evaluating error detection and repair models, as the synthetic errors may not align with the challenges encountered in real-world applications.
@@ 2.2.3. Leveraging Instruction-Tuned LLMs for Table Error Processing
To address these limitations, recent studies have explored instruction-tuned LLMs that explicitly integrate table-specific constraints and domain knowledge into their fine-tuning process (li2024table). Our approach builds on this paradigm by incorporating structural metadata—such as schema constraints and functional dependencies—into the LLM’s training pipeline. By embedding such constraints within the instruction design, we guide the model to generate contextually valid errors that better reflect real-world data corruptions while maintaining table consistency.
@@ 3.3. Datasets
By ensuring a balanced combination of these factors, we aim to build a model capable of generating realistic and diverse errors applicable to a broad range of tabular data. The datasets used in this study are primarily sourced from those introduced in (li2021cleanml; mahdavi2019raha), which are primarily designed for error detection and data cleaning tasks.
@@
    The Error-unrelated tasks are inspired by Table-GPT (li2024table) and are designed to enhance the model’s ability to recognize row and column positions while improving its overall understanding of tabular data.
@@ 4.1. Table Task Summary
In Section 1, we discuss that our primary objective is to enhance the ability of LLMs to generate realistic errors. To this end, we carefully select a set of tasks for fine-tuning the model. Drawing inspiration from previous works (li2024table; zhang2023tablellama), we first incorporate a series of error-agnostic tasks to bolster the basic reading and writing capabilities of LLMs on tabular data. These tasks enable the model to analyze two-dimensional tables and accurately identify specific rows and columns. For instance, as illustrated in Table 2, Row/Column Augmentation (Task-4) improves the model’s understasnding of row and column identifiers, Header Matching (Task-7), Row/Column Swapping (Task-5), and Filtering (Task-6) enhance its ability to read positional information, and Table Summarization (Task-8) reinforces its comprehension of table content.
@@ 4.2. Prompt Builder
To address the aforementioned challenges in table-tuning and error-related tasks, we build on instruction fine-tuning (ouyang2022training) and the table-tuning paradigm introduced by Li et al. (li2024table). Specifically, we design a triplet representation $(I,T,O)$ to tackle error-related tasks in tabular data.
@@ 4.3. TableEG Training Procedure
We fine-tune (brown2020language) our TableEG model using the LLaMA3.1-8B base model with LoRA-based adaptation. The training procedure is organized as follows:
@@ 4.3.1. Model and Adaptation:
We employ the LLaMA3.1-8B model as the base, and apply LoRA (hu2022lora) to its projection layers (i.e., $\{q\_proj,k\_proj,v\_proj,o\_proj\}$) to efficiently adapt the model for table-related error generation tasks.
@@ 4.3.3. Training Data and Tasks:
We construct the training set from error-labeled datasets (refer to Section 3.3) with a 90:10 split (90% for training). The training tasks include error generation, error detection, and error correction. Additionally, we incorporate synthetic tasks from Table-GPT (li2024table) (e.g., column/row augmentation, swapping, filtering, sorting, table summarization, header matching, and natural language to SQL conversion) to enhance the model’s understanding of two-dimensional table structures.
@@ 7. Related Work
Ensuring data quality in relational databases and tabular data processing remains a significant challenge, as errors like missing values, pattern violations, and rule violations can disrupt downstream analytical and machine learning tasks. Traditional rule-based approaches, such as functional dependencies (FDs), uniqueness constraints (UCs), and pattern constraints, have been widely used (rekatsinas2017holoclean). However, these methods require extensive manual specification, limiting their adaptability to large-scale, high-dimensional data. Recent advancements in deep learning and large language models (LLMs) have introduced more flexible solutions for data quality management. Transformer-based models have shown strong generalization across various tasks, including data imputation, error detection, and error correction (li2024table; deng2022turl). AI-driven methods further enhance adaptability by learning complex data patterns and reducing manual intervention (zhu2024relational).
@@ 7.1. Data Imputation
Data imputation is essential for maintaining data completeness. Traditional methods include statistical techniques (e.g., mean, median, mode (wei2018missing)), matrix factorization (SVD (troyanskaya2001missing), PCA (oba2003bayesian)), regression-based approaches (LOESS (cleveland1996smoothing), IIM (zhang2019learning)), and nearest-neighbor imputation (kNN (regression1992introduction), MIBOS (wu2012missing)). Constraint-based methods leverage data integrity constraints (MLEM2 (grzymala2005handling), DERAND (song2015enriching; song2018enriching)), while multiple imputation techniques (MICE (van2011mice), SICE (khan2020sice)) generate multiple candidate values to improve robustness. However, these approaches often rely on strong assumptions, limiting their adaptability to complex data structures.
Machine learning (ML) and deep learning (DL) have improved imputation by capturing intricate data dependencies. Neural network-based models like RNNI (rumelhart1986learning) and MLPI (garcia2010pattern) enhance generalization, while autoencoders (MIDAE (gondara2018mida), SDAE (costa2018missing)), VAEs (HI-VAE (nazabal2020handling)), and graph neural networks (GRAPE (you2020handling), IGRM (zhong2023data)) provide structural awareness. GAN-based approaches (GAIN (yoon2018gain), WGAN (arjovsky2017wasserstein)) further refine imputation by generating plausible missing values. More recently, large language models (LLMs) like GPT-3 (narayan2022can) and Table-GPT (li2024table) leverage zero-shot and few-shot learning for imputation, offering strong generalization but facing limitations in handling numerical data. Despite these advancements, DL-based methods remain computationally expensive and less interpretable than traditional approaches.
@@ 7.2. Error Detection
Error detection is essential for identifying anomalies in relational data. Traditional methods include statistical techniques (e.g., PCA (shyu2003novel), dBoost (pit2016outlier)) and constraint-based approaches enforcing integrity constraints (e.g., FDs (mandros2017discovering), CFDs (fan2008conditional), DCs (chu2013discovering)). Additionally, clustering-based methods (e.g., DBSCAN (ester1996density), OPTICS (ankerst1999optics)) and outlier detection techniques (e.g., LOF (breunig2000lof), iForest (liu2008isolation)) have also been widely used to capture anomalies.
Beyond traditional approaches, machine learning (ML)-based approaches enhance detection by integrating multiple detection signals and classifier models (e.g., Raha (mahdavi2019raha), HoloDetect (heidari2019holodetect)). Unsupervised (e.g., Uni-Detect (wang2019uni)) and active learning techniques (e.g., ED2 (neutatz2019ed2)) further improve adaptability by reducing reliance on labeled data. Deep learning (DL)-based methods advance error detection by employing autoencoders for anomaly detection (e.g., RandNet (chen2017outlier), REPEN (pang2019deep)) and utilizing semi-supervised learning to model normal and erroneous data distributions (e.g., DeepSAD (ruff2019deep)). Recently, large language models (LLMs) have been fine-tuned for tabular data anomaly detection (e.g., ADTDL (li2024anomaly)), showing potential despite computational challenges.

### v3

@@ 1. Introduction
Ensuring high data quality is a fundamental challenge in data-driven systems, as erroneous tabular data can terribly degrade the reliability of analytical processes and machine learning models (Rahm et al., 2000; Chu et al., 2016). To tackle these errors, researchers spend a lot of time designing various data cleaning techniques (Chu et al., 2013b; Mahdavi et al., 2019; Rezig et al., 2021). As we know, sufficient training data with labeled errors are important for the performance of data cleaning models (Rekatsinas et al., 2017). Unfortunately, the manual labeling is very expensive, and obtaining a sufficiently diverse set of real-world errors remains challenging (Krishnan et al., 2017). In this sense, it is natural to come up with the idea that we can manually generate errors for the model training.
Existing error generation methods primarily rely on rule-based approaches, with BART (Arocena et al., 2015) being the most representative framework. BART introduces errors into clean datasets by applying predefined constraints, such as functional dependencies (FDs) (Mandros et al., 2017; Fan et al., 2008) and denial constraints (DCs) (Bian et al., 2024; Pena et al., 2022; Chu et al., 2013a), ensuring that the generated errors are systematically detectable. By enforcing a controlled corruption process, BART provides a benchmark for evaluating data cleaning algorithms.
Figure 1. An example of the generated errors by BART (Arocena et al., 2015) and our TableEG, over the Movie dataset.
@@
Second, BART focuses predominantly on constraint-violating errors. Although BART is designed to support a variety of random errors, including typos, duplicated values, outliers, and even missing or bogus values, its rule-based mechanism is inherently limited. In practice, BART struggles to generate missing values (Krishnan et al., 2017) that faithfully reflect the nuanced patterns present in real-world datasets and to produce domain-specific semantic inconsistencies (Suchanek et al., 2011). For instance, in the Flights (Rekatsinas et al., 2017) dataset, the attribute act_arr_time may contain the value “Not available,” indicating missing or ambiguous data rather than a simple constraint violation. However, BART cannot generate such errors as it assumes all errors stem from various constraint violations. This limitation also weakens its suitability for benchmarking error detection algorithms, which must handle a broader and more diverse range of real-world data corruptions.
@@
Recent advancements in Large Language Models (LLMs) have achieved remarkable success in natural language processing (NLP) tasks, drawing increasing attention to their potential for structured tabular data processing tasks (Li et al., 2024b; Wu et al., 2024; Zhang et al., 2023). In this sense, it is intuitive to leverage LLMs’ robust semantic comprehension and expansive knowledge to systematically generate authentic errors. However, unlike natural language, relational data requires maintaining two-dimensional structural integrity while accounting for inter-cell dependencies (e.g., primary-foreign key relationships) and domain-specific constraints (e.g., financial transaction rules). This raises the fundamental questions: Can LLMs generate contextually meaningful errors rather than entirely random or nonsensical corruptions? Do these synthetic errors faithfully replicate the diverse real-world errors so that they can be reliably detected by current error detection algorithms?
@@ 2.1. Types of Data Errors
In data cleaning and quality research, errors are typically defined as deviations of data values from their ground truth (Abedjan et al., 2016).
@@
These errors can arise from various reasons—such as noise in data acquisition, manual entry mistakes, rule violations, or inconsistencies introduced during data integration. Several studies have proposed classification schemes to categorize these errors either based on their origins (Rahm et al., 2000; Kim et al., 2003; Hellerstein, 2013) or on the violations of constraints and patterns observed in structured data (Ilyas et al., 2015).
Our classification scheme is broadly consistent with the taxonomy proposed by Abedjan et al. (Abedjan et al., 2016), which, although not exhaustive, serves as a well-established framework for systematically categorizing errors commonly identified in real-world datasets. However, our classification differs in several key aspects. First, we explicitly separate missing values as an independent category, acknowledging their prevalence in real-world data. Second, we incorporate duplicates within the broader category of rule violations. Lastly, our definition of rule violations extends to any value that violates a form of denial constraints, encompassing a wider range of integrity violations. Consequently, we categorize errors into four primary types—outliers, missing values, rule violations, and pattern violations—capturing the most prevalent and impactful error types observed in real-world tabular datasets.
@@
Rule Violations: We employ a set of predefined integrity constraints, primarily in the form of DCs (Chu et al., 2013a)), FDs (Mandros et al., 2017) and CFDs (Fan et al., 2008) to safeguard data consistency. Any record that fails to meet these constraints is classified as a rule violation. Such violations often involve contradictory, prohibited, or otherwise infeasible values, thereby severely undermining data integrity. They typically arise from data entry errors, integration inconsistencies, or improper transformations, and are commonly detected through constraint validation or rule-based inference.
@@ 2.2.1. Recent Advancements in Large Language Models (LLMs)
Recent advancements in Large Language Models (LLMs) have demonstrated remarkable capabilities in natural language understanding and generation, primarily trained on vast corpora of sequential text (Vaswani et al., 2017; Brown et al., 2020). However, directly applying LLMs to relational data and tabular tasks remains nontrivial due to the inherent structural differences between natural language and relational tables. Unlike textual data, which follows a linear sequence, tabular data is inherently two-dimensional, often governed by schema constraints, cross-column dependencies, and entity relationships (Dong and Wang, 2024). These structural properties impose significant challenges when leveraging LLMs for table-based tasks such as error generation, detection, and repair.
@@ 2.2.2. Challenges of Applying LLMs to Tabular Data
First, traditional LLMs primarily operate under a token-based autoregressive or masked prediction framework, which excels in modeling linguistic dependencies but struggles with structured tabular constraints (Yin et al., 2020). In relational tables, data integrity often depends on multi-column dependencies, such as functional dependencies (FDs) or domain constraints, which LLMs do not inherently recognize (Chu et al., 2016). Consequently, naive applications of LLMs to table-based tasks often result in syntactically plausible yet semantically invalid outputs, where column dependencies are violated, inconsistent values are introduced, or key constraints are broken. Second, LLMs tend to generate errors based on linguistic priors rather than data-specific corruption patterns. Unlike real-world data errors—which may arise from user input mistakes, OCR misrecognition, or systematic extraction faults—LLM-generated corruptions often exhibit artificial substitution biases that fail to reflect natural error distributions. This discrepancy limits the effectiveness of LLM-generated datasets for evaluating error detection and repair models, as the synthetic errors may not align with the challenges encountered in real-world applications.
@@ 2.2.3. Leveraging Instruction-Tuned LLMs for Table Error Processing
To address these limitations, recent studies have explored instruction-tuned LLMs that explicitly integrate table-specific constraints and domain knowledge into their fine-tuning process (Li et al., 2024b). Our approach builds on this paradigm by incorporating structural metadata—such as schema constraints and functional dependencies—into the LLM’s training pipeline. By embedding such constraints within the instruction design, we guide the model to generate contextually valid errors that better reflect real-world data corruptions while maintaining table consistency.
@@ 3.3. Datasets
By ensuring a balanced combination of these factors, we aim to build a model capable of generating realistic and diverse errors applicable to a broad range of tabular data. The datasets used in this study are primarily sourced from those introduced in (Li et al., 2021; Mahdavi et al., 2019), which are primarily designed for error detection and data cleaning tasks.
@@
    The Error-unrelated tasks are inspired by Table-GPT (Li et al., 2024b) and are designed to enhance the model’s ability to recognize row and column positions while improving its overall understanding of tabular data.
@@ 4.1. Table Task Summary
In Section 1, we discuss that our primary objective is to enhance the ability of LLMs to generate realistic errors. To this end, we carefully select a set of tasks for fine-tuning the model. Drawing inspiration from previous works (Li et al., 2024b; Zhang et al., 2023), we first incorporate a series of error-agnostic tasks to bolster the basic reading and writing capabilities of LLMs on tabular data. These tasks enable the model to analyze two-dimensional tables and accurately identify specific rows and columns. For instance, as illustrated in Table 2, Row/Column Augmentation (Task-4) improves the model’s understasnding of row and column identifiers, Header Matching (Task-7), Row/Column Swapping (Task-5), and Filtering (Task-6) enhance its ability to read positional information, and Table Summarization (Task-8) reinforces its comprehension of table content.
@@ 4.2. Prompt Builder
To address the aforementioned challenges in table-tuning and error-related tasks, we build on instruction fine-tuning (Ouyang et al., 2022) and the table-tuning paradigm introduced by Li et al. (Li et al., 2024b). Specifically, we design a triplet representation $(I,T,O)$ to tackle error-related tasks in tabular data.
@@ 4.3. TableEG Training Procedure
We fine-tune (Brown et al., 2020) our TableEG model using the LLaMA3.1-8B base model with LoRA-based adaptation. The training procedure is organized as follows:
@@ 4.3.1. Model and Adaptation:
We employ the LLaMA3.1-8B model as the base, and apply LoRA (Hu et al., 2022) to its projection layers (i.e., $\{q\_proj,k\_proj,v\_proj,o\_proj\}$) to efficiently adapt the model for table-related error generation tasks.
@@ 4.3.3. Training Data and Tasks:
We construct the training set from error-labeled datasets (refer to Section 3.3) with a 90:10 split (90% for training). The training tasks include error generation, error detection, and error correction. Additionally, we incorporate synthetic tasks from Table-GPT (Li et al., 2024b) (e.g., column/row augmentation, swapping, filtering, sorting, table summarization, header matching, and natural language to SQL conversion) to enhance the model’s understanding of two-dimensional table structures.
@@ 7. Related Work
Ensuring data quality in relational databases and tabular data processing remains a significant challenge, as errors like missing values, pattern violations, and rule violations can disrupt downstream analytical and machine learning tasks. Traditional rule-based approaches, such as functional dependencies (FDs), uniqueness constraints (UCs), and pattern constraints, have been widely used (Rekatsinas et al., 2017). However, these methods require extensive manual specification, limiting their adaptability to large-scale, high-dimensional data. Recent advancements in deep learning and large language models (LLMs) have introduced more flexible solutions for data quality management. Transformer-based models have shown strong generalization across various tasks, including data imputation, error detection, and error correction (Li et al., 2024b; Deng et al., 2022). AI-driven methods further enhance adaptability by learning complex data patterns and reducing manual intervention (Zhu et al., 2024).
@@ 7.1. Data Imputation
Data imputation is essential for maintaining data completeness. Traditional methods include statistical techniques (e.g., mean, median, mode (Wei et al., 2018)), matrix factorization (SVD (Troyanskaya et al., 2001), PCA (Oba et al., 2003)), regression-based approaches (LOESS (Cleveland and Loader, 1996), IIM (Zhang et al., 2019)), and nearest-neighbor imputation (kNN (Regression, 1992), MIBOS (Wu et al., 2012)). Constraint-based methods leverage data integrity constraints (MLEM2 (Grzymala-Busse et al., 2005), DERAND (Song et al., 2015b, 2018)), while multiple imputation techniques (MICE (Van Buuren and Groothuis-Oudshoorn, 2011), SICE (Khan and Hoque, 2020)) generate multiple candidate values to improve robustness. However, these approaches often rely on strong assumptions, limiting their adaptability to complex data structures.
Machine learning (ML) and deep learning (DL) have improved imputation by capturing intricate data dependencies. Neural network-based models like RNNI (Rumelhart et al., 1986) and MLPI (García-Laencina et al., 2010) enhance generalization, while autoencoders (MIDAE (Gondara and Wang, 2018), SDAE (Costa et al., 2018)), VAEs (HI-VAE (Nazabal et al., 2020)), and graph neural networks (GRAPE (You et al., 2020), IGRM (Zhong et al., 2023)) provide structural awareness. GAN-based approaches (GAIN (Yoon et al., 2018), WGAN (Arjovsky et al., 2017)) further refine imputation by generating plausible missing values. More recently, large language models (LLMs) like GPT-3 (Narayan et al., 2022) and Table-GPT (Li et al., 2024b) leverage zero-shot and few-shot learning for imputation, offering strong generalization but facing limitations in handling numerical data. Despite these advancements, DL-based methods remain computationally expensive and less interpretable than traditional approaches.
@@ 7.2. Error Detection
Error detection is essential for identifying anomalies in relational data. Traditional methods include statistical techniques (e.g., PCA (Shyu et al., 2003), dBoost (Pit-Claudel et al., 2016)) and constraint-based approaches enforcing integrity constraints (e.g., FDs (Mandros et al., 2017), CFDs (Fan et al., 2008), DCs (Chu et al., 2013a)). Additionally, clustering-based methods (e.g., DBSCAN (Ester et al., 1996), OPTICS (Ankerst et al., 1999)) and outlier detection techniques (e.g., LOF (Breunig et al., 2000), iForest (Liu et al., 2008)) have also been widely used to capture anomalies.
Beyond traditional approaches, machine learning (ML)-based approaches enhance detection by integrating multiple detection signals and classifier models (e.g., Raha (Mahdavi et al., 2019), HoloDetect (Heidari et al., 2019)). Unsupervised (e.g., Uni-Detect (Wang and He, 2019)) and active learning techniques (e.g., ED2 (Neutatz et al., 2019)) further improve adaptability by reducing reliance on labeled data. Deep learning (DL)-based methods advance error detection by employing autoencoders for anomaly detection (e.g., RandNet (Chen et al., 2017), REPEN (Pang et al., 2019)) and utilizing semi-supervised learning to model normal and erroneous data distributions (e.g., DeepSAD (Ruff et al., 2019)). Recently, large language models (LLMs) have been fine-tuned for tabular data anomaly detection (e.g., ADTDL (Li et al., 2024a)), showing potential despite computational challenges.